# Speech-Aware Long Context Pruning and Integration for Contextualized Automatic Speech Recognition

### v1

@@ 1 Introduction
Automatic speech recognition (ASR) systems have achieved remarkable performance in common scenarios, approaching human-level accuracy in routine speech recognition tasks (radford2023robust; gao2022paraformer; xu2025fireredasr). However, current challenges no longer center around baseline transcription accuracy but instead focus on effectively recognizing speech in complex, context-rich scenarios (chang2021context; ni2024vilas; wang2024cieasr; sudo2024contextualized). For example, in conference videos, speakers commonly utilize visual aids, particularly slides containing large amounts of text (wang2024slidespeech; wang2024slideavsr). Such OCR-derived contexts usually include important entities and domain-specific terminology. In Figure 1, despite that keywords like ”glaucoma” constitute only a small portion of the speech content, they are crucial for comprehending domain-specific information. However, processing extensive textual inputs introduces considerable computational overhead and the risk of incorporating irrelevant information when the textual contexts span hundreds of thousands of words.
Existing contextualized ASR methods mainly follow two paradigms: (1) deep biasing E2E models that apply attention mechanisms to influence the decoding of ASR model (pundak2018deep; chang2021context; han2021cif; han2022improving; huang2023contextualized; sudo2024contextualized; yu2024lcb), and (2) SpeechLLM models that treat contextual keywords as prompts (lakomkin2024end; yang2024mala; bai2024seed). Deep biasing models face performance gaps compared with large-scale pre-trained models (lakomkin2024end) and exhibit architectural rigidity. SpeechLLM models are more flexible and robust when incorporating contextual keywords. However, when dealing with long context information, simply treating contextual keywords as prompts poses challenges to effective long-context processing and robust noise filtering.
Neuroscience study on speech recognition (sohoglu2012predictive) indicates that human perceptual clarity improves significantly when speech aligns with prior contextual knowledge, demonstrating the human brain’s active integration of prior knowledge during speech recognition. Our motivation is to simulate this active integration process, prune and compress noisy contextual information to knowledge more aligned with speech contents. We propose SAP2 (Speech-Aware Context Pruning with Speech-Driven Attention-based Pooling), a novel framework that actively filters long-form contextual keywords and integrates them into ASR systems without sacrificing efficiency. Our contributions are as follows:
@@ 2.2 Speech-Aware Context Pruning
This derivation decomposes the original task into two stages. Prior work like jayanthi2023retrieve explores this two-stage prune-and-integrate pipeline. However, the pruning model $p_{\psi}$ is mainly based on similarity-based algorithms, which makes it hard to generalize to unseeing textual contents and speech. In Figure 2, we utilize a SpeechLLM model as the context pruning model. It accepts the input of long contextual keywords and speech, and output core keywords relevant to the speech using speech-driven attention-based pooling. The core keywords generated in step one is concatenated with the speech to perform contextualized ASR. Regarding this Two-stage Pruning-Integration (TPI) manager, we denote the pipeline as SAP2-TPI. Unless specified otherwise, SAP2 refers to the two-phase trained variant (SAP2-TPI). During the training time, ASR model $p_{\theta}$ and speech-driven context pruning model $p_{\psi}$ can be trained simultaneously, which means the training objective can be written as
@@ 3.1 Experimental Setup
We conduct experiments on the SlideSpeech (wang2024slidespeech) dataset and the LibriSpeech (LibriSpeech) dataset. The contextual keywords of SlideSpeech are slide-derived OCR texts, and the contextual keywords of LibriSpeech is dynamically constructed following the approach in Le2021ContextualizedSE-DB.
@@
We follow the evaluation metrics of previous works of contextualized ASR (huang2023contextualized; yu2024lcb; yang2024mala; wang2024slidespeech). We evaluate the results using WER, biased word error rate (B-WER), unbiased word error rate (U-WER), and the Recall of words in both the biasing list and transcript. U-WER is calculated for words not in the contextual keywords list, which reflects the foundational ASR capabilities of the model. B-WER is computed for words in the contextual keywords list, which indicates the model’s capability of accurately integrating core keywords. For contextualized ASR models, U-WER is mainly related to foundation ASR models, so it’s more important to achieve improvements in B-WER while maintaining U-WER stability.
@@
Train Model
Contextual
Keywords
Test
@@
S95
SlidesSpeech (wang2024slidespeech) 21.22 20.83 26.60 73.51
MaLa-ASR (yang2024mala) 11.80 11.71 13.52 86.71
Qwen2-Audio (chu2024qwen2) 10.79 10.85 7.92 90.33
CPP (huang2023contextualized) 1slide 20.95 20.73 24.05 76.10
LCB-net (yu2024lcb) 1slide 19.21 18.89 23.70 76.48
MaLa-ASR (yang2024mala) 1slide 11.26 11.52 7.67 92.50
@@
L95
SlidesSpeech (wang2024slidespeech) 12.89 12.90 12.70 87.43
MaLa-ASR (yang2024mala) 8.61 8.72 7.34 92.84
Qwen2-Audio (chu2024qwen2) 10.76 10.91 7.71 91.42
CPP (huang2023contextualized) 1slide 12.38 12.60 9.32 90.86
LCB-net (yu2024lcb) 1slide 12.02 12.24 9.03 91.12
MaLa-ASR (yang2024mala) 1slide 8.46 8.73 4.89 95.31
@@ 3.2 Implementation Details
DB-RNNT+DB-LM (Le2021ContextualizedSE-DB)
1.98
(1.5/5.7)
5.86
(4.9/14.1)
2.09
(1.6/6.2)
6.09
(5.1/15.1)
2.14
(1.6/6.7)
6.35
(5.1/17.2)
Attention-based DB
+ BPB beam search (sudo2024contextualized)
2.75
(2.3/6.0)
5.60
(4.9/12.0)
3.21
(2.7/7.0)
6.28
(5.5/13.5)
3.47
(3.0/7.7)
7.34
(6.4/15.8)
Biasing fusion
+ SpeechLLM (gong24b_biasfusion+speechllm)
1.6
(1.3/5.5)
3.8
(2.6/13.5)
1.9
(1.4/6.0)
3.9
(2.7/14.2)
CTC-Assisted LLM-Based ASR
1.27
(1.00/3.67)
2.72
(2.16/8.02)
1.33
(1.03/3.92)
3.04
(2.40/9.04)
1.33
(1.00/4.16)
2.99
(2.31/9.33)
SAP2-TPI (Proposed)
1.12
(0.89/3.03)
2.60
(2.18/6.55)
1.16
(0.93/3.09)
2.72
(2.23/7.29)
1.22
(0.96/3.50)
2.94
(2.33/8.62)
@@ 3.4 Experiment Results on LibriSpeech
  • SAP2-TPI also achieves a SOTA performance on LibriSpeech across all biasing list settings, reaching an lowest WER of 1.12%. SAP2-TPI achieves absolute reductions of 0.48% and 0.74% on WER compared with previous SOTA gong24b_biasfusion+speechllm on test-clean (N=100) and test-clean (N=500), with even large absolute reductions of 0.8% and 1.18% on test-other (N=100) and test-other(N=100).
@@ 3.5 Ablation Studies
Prompt Concatenation (PC) Following yang2024mala, we directly concatenate contextual keywords with speech features. We design different instructions according to two situations: contextual keywords are provided, or no keywords are available. Please refer to the Appendix for detailed prompts.
@@
Train Model
Contextual
Keywords
Test
@@
Model
Contextual
Keywords
F1-score
@@
Train
Contextual
Keywords
Model Test
@@
L95
1slide Qwen2-Audio
@@
Method
Training
time (s)
Inference time (s)
(15 slides)
@@
Qwen2-Audio-TPI
(Stage I )
51003.8 1960
Qwen2-Audio-TPI
(Stage II )
30120.3 5047
@@ A.1 Pre-trained Speech Models
The field of speech recognition has made remarkable strides with unsupervised pre-training techniques like Wav2Vec 2.0  (baevski2020wav2vec) and w2v-BERT (chung2021w2v). A key advancement in this area is Whisper (radford2023robust), which is pre-trained in a weakly supervised manner on large-scale speech datasets. Its robust pre-training capabilities, coupled with a flexible Encoder-Decoder architecture, have made it a popular choice for a variety of downstream tasks (wang2024can; wang2024cieasr; attia2024kid). The emergence of large multimodal models has driven further progress, particularly for tasks that require the alignment of speech and language. Models like Seed-ASR (bai2024seed), Qwen2-Audio (chu2024qwen2), and FireRedASR (xu2025fireredasr) offer greater flexibility and precision across diverse speech-related tasks.
@@ A.2 Context Compression
Recent approaches to context compression in large language models (LLMs) focus on improving the efficiency of handling long contexts while retaining important information. chevalier2023adapting compresses long contexts into summary vectors, which are then used as soft prompts to improve model performance without significantly increasing the computational cost. ge2023context utilizes a pre-trained autoencoder to generate compact memory slots for context. li2023compressing proposes a method that prunes redundant information in the input context, effectively reducing its size while maintaining key contextual elements. LongLLMLingua (jiang2023llmlingua; jiang2023longllmlingua) focuses on prompt compression for long-context scenarios, using a mechanism to enhance the LLM’s ability to retain key information in the compressed prompt. cao2024retaining uses queries to guide the compression process, ensuring important contextual information is retained during compression even with high compression ratios.
@@ A.3 Contextualized ASR
Recent advances in Contextualized ASR have shown significant improvements by incorporating contextual information. A common approach is deep biasing end-to-end models. huang2023contextualized utilizes textual context phrases to assist in training ASR models, improving performance in recognizing longer, context-dependent sequences. sudo2024contextualized uses an editable bias phrase list to inject domain-specific knowledge into the ASR system. yang2024promptasr incorporates text prompts into end-to-end ASR systems, enabling flexible transcription control and contextual understanding. jayanthi2023retrieve explores a two-stage retrieve-and-copy pipeline to retrieve related textual content based on similarity metrics. The other approaches utilize SpeechLLM models for enhanced contextual modeling capabilities. lakomkin2024end uses LLaMA for speech recognition, treating the task as a mixed-modal language modeling problem and effectively incorporating textual contexts alongside audio inputs. yang2024mala utilizes the WavLM Large model as a speech encoder, pre-trained on extensive audio data, and Vicuna as its LLM decoder.
@@ C.1 Dataset
We conduct experiments on the SlideSpeech (wang2024slidespeech) dataset. This large-scale multimodal collection contains 1,705 YouTube-sourced conference videos, featuring synchronized 720p video streams, 16kHz audio recordings, and slide-derived OCR texts. The corpus uniquely integrates pre-processed slide content including extracted keywords and layout annotations. The training set is stratified into two subsets: L95 and S95, comprising 473 hours and 161 hours of audio data, respectively.
@@
To more fairly and comprehensively evaluate the performance of our proposed SAP2, we conduct experiments on the LibriSpeech (LibriSpeech) corpus. For the training and validation sets, we follow the approach in Le2021ContextualizedSE-DB to dynamically construct biasing lists. Specifically, for each sample, we first identify words in its ASR label that fall outside the common5k vocabulary. Then, we add $N$ (randomly selected between 400 and 800) distractors randomly sampled from the pool of 209.2k rare words. The test set remains consistent with Le2021ContextualizedSE-DB. In line with most previous works, we evaluate our method under three settings: $N=\{100,500,1000\}$. We train separately on the 460-hour ”train-clean” set and the 500-hour ”train-other” set, and perform inference on the ”test-clean” and ”test-other” sets, respectively.

### v2

@@ 1 Introduction
Automatic speech recognition (ASR) systems have achieved remarkable performance in common scenarios, approaching human-level accuracy in routine speech recognition tasks (Radford et al. 2023; Gao et al. 2022; Xu et al. 2025). However, current challenges no longer center around baseline transcription accuracy but instead focus on effectively recognizing speech in complex, context-rich scenarios (Chang et al. 2021; Ni et al. 2024; Wang et al. 2024d; Sudo et al. 2024). For example, in conference videos, speakers commonly utilize visual aids, particularly slides containing large amounts of text (Wang et al. 2024b, a). Such OCR-derived contexts usually include important entities and domain-specific terminology. In Figure 1, despite that keywords like ”glaucoma” constitute only a small portion of the speech content, they are crucial for comprehending domain-specific information. However, processing extensive textual inputs introduces considerable computational overhead and the risk of incorporating irrelevant information when the textual contexts span hundreds of thousands of words.
Existing contextualized ASR methods mainly follow two paradigms: (1) deep biasing E2E models that apply attention mechanisms to influence the decoding of ASR model (Pundak et al. 2018; Chang et al. 2021; Han et al. 2021, 2022; Huang et al. 2023; Sudo et al. 2024; Yu et al. 2024), and (2) SpeechLLM models that treat contextual keywords as prompts (Lakomkin et al. 2024; Yang et al. 2024b; Bai et al. 2024). Deep biasing models face performance gaps compared with large-scale pre-trained models (Lakomkin et al. 2024) and exhibit architectural rigidity. SpeechLLM models are more flexible and robust when incorporating contextual keywords. However, when dealing with long context information, simply treating contextual keywords as prompts poses challenges to effective long-context processing and robust noise filtering.
Neuroscience study on speech recognition (Sohoglu et al. 2012) indicates that human perceptual clarity improves significantly when speech aligns with prior contextual knowledge, demonstrating the human brain’s active integration of prior knowledge during speech recognition. Our motivation is to simulate this active integration process, prune and compress noisy contextual information to knowledge more aligned with speech contents. We propose SAP2 (Speech-Aware Context Pruning with Speech-Driven Attention-based Pooling), a novel framework that actively filters long-form contextual keywords and integrates them into ASR systems without sacrificing efficiency. Our contributions are as follows:
@@ 2.2 Speech-Aware Context Pruning
This derivation decomposes the original task into two stages. Prior work like Jayanthi et al. (2023) explores this two-stage prune-and-integrate pipeline. However, the pruning model $p_{\psi}$ is mainly based on similarity-based algorithms, which makes it hard to generalize to unseeing textual contents and speech. In Figure 2, we utilize a SpeechLLM model as the context pruning model. It accepts the input of long contextual keywords and speech, and output core keywords relevant to the speech using speech-driven attention-based pooling. The core keywords generated in step one is concatenated with the speech to perform contextualized ASR. Regarding this Two-stage Pruning-Integration (TPI) manager, we denote the pipeline as SAP2-TPI. Unless specified otherwise, SAP2 refers to the two-phase trained variant (SAP2-TPI). During the training time, ASR model $p_{\theta}$ and speech-driven context pruning model $p_{\psi}$ can be trained simultaneously, which means the training objective can be written as
@@ 3.1 Experimental Setup
We conduct experiments on the SlideSpeech (Wang et al. 2024b) dataset and the LibriSpeech (Panayotov et al. 2015) dataset. The contextual keywords of SlideSpeech are slide-derived OCR texts, and the contextual keywords of LibriSpeech is dynamically constructed following the approach in Le et al. (2021).
@@
We follow the evaluation metrics of previous works of contextualized ASR (Huang et al. 2023; Yu et al. 2024; Yang et al. 2024b; Wang et al. 2024b). We evaluate the results using WER, biased word error rate (B-WER), unbiased word error rate (U-WER), and the Recall of words in both the biasing list and transcript. U-WER is calculated for words not in the contextual keywords list, which reflects the foundational ASR capabilities of the model. B-WER is computed for words in the contextual keywords list, which indicates the model’s capability of accurately integrating core keywords. For contextualized ASR models, U-WER is mainly related to foundation ASR models, so it’s more important to achieve improvements in B-WER while maintaining U-WER stability.
@@
Train Model Contextual Keywords Test
@@
S95 (161h) SlidesSpeech (Wang et al. 2024b) 21.22 20.83 26.60 73.51
MaLa-ASR (Yang et al. 2024b) 11.80 11.71 13.52 86.71
Qwen2-Audio (Chu et al. 2024) 10.79 10.85 7.92 90.33
CPP (Huang et al. 2023) 1slide 20.95 20.73 24.05 76.10
LCB-net (Yu et al. 2024) 1slide 19.21 18.89 23.70 76.48
MaLa-ASR (Yang et al. 2024b) 1slide 11.26 11.52 7.67 92.50
@@
L95 (473h) SlidesSpeech (Wang et al. 2024b) 12.89 12.90 12.70 87.43
MaLa-ASR (Yang et al. 2024b) 8.61 8.72 7.34 92.84
Qwen2-Audio (Chu et al. 2024) 10.76 10.91 7.71 91.42
CPP (Huang et al. 2023) 1slide 12.38 12.60 9.32 90.86
LCB-net (Yu et al. 2024) 1slide 12.02 12.24 9.03 91.12
MaLa-ASR (Yang et al. 2024b) 1slide 8.46 8.73 4.89 95.31
@@ 3.2 Implementation Details
DB-RNNT+DB-LM (Le et al. 2021) 1.98 (1.5/5.7) 5.86 (4.9/14.1) 2.09 (1.6/6.2) 6.09 (5.1/15.1) 2.14 (1.6/6.7) 6.35 (5.1/17.2)
Attention-based DB + BPB beam search (Sudo et al. 2024) 2.75 (2.3/6.0) 5.60 (4.9/12.0) 3.21 (2.7/7.0) 6.28 (5.5/13.5) 3.47 (3.0/7.7) 7.34 (6.4/15.8)
Biasing fusion + SpeechLLM (Gong et al. 2024) 1.6 (1.3/5.5) 3.8 (2.6/13.5) 1.9 (1.4/6.0) 3.9 (2.7/14.2)
CTC-Assisted LLM-Based ASR (Yang et al. 2024a) 1.27 (1.00/3.67) 2.72 (2.16/8.02) 1.33 (1.03/3.92) 3.04 (2.40/9.04) 1.33 (1.00/4.16) 2.99 (2.31/9.33)
SAP2-TPI (Proposed) 1.12 (0.89/3.03) 2.60 (2.18/6.55) 1.16 (0.93/3.09) 2.72 (2.23/7.29) 1.22 (0.96/3.50) 2.94 (2.33/8.62)
@@ 3.4 Experiment Results on LibriSpeech
  • SAP2-TPI also achieves a SOTA performance on LibriSpeech across all biasing list settings, reaching an lowest WER of 1.12%. SAP2-TPI achieves absolute reductions of 0.48% and 0.74% on WER compared with previous SOTA Gong et al. (2024) on test-clean (N=100) and test-clean (N=500), with even large absolute reductions of 0.8% and 1.18% on test-other (N=100) and test-other(N=100).
@@ 3.5 Ablation Studies
Prompt Concatenation (PC) Following Yang et al. (2024b), we directly concatenate contextual keywords with speech features. We design different instructions according to two situations: contextual keywords are provided, or no keywords are available. Please refer to the Appendix for detailed prompts.
@@
Train Model Contextual Keywords Test
@@
Model Contextual Keywords F1-score
@@
Train Contextual Keywords Model Test
@@
L95 (473h) 1slide Qwen2-Audio
@@
Method Training time (s) Inference time (s) (15 slides)
@@
Qwen2-Audio-TPI (Stage I ) 51003.8 1960
Qwen2-Audio-TPI (Stage II ) 30120.3 5047
@@ A.1 Pre-trained Speech Models
The field of speech recognition has made remarkable strides with unsupervised pre-training techniques like Wav2Vec 2.0  (Baevski et al. 2020) and w2v-BERT (Chung et al. 2021). A key advancement in this area is Whisper (Radford et al. 2023), which is pre-trained in a weakly supervised manner on large-scale speech datasets. Its robust pre-training capabilities, coupled with a flexible Encoder-Decoder architecture, have made it a popular choice for a variety of downstream tasks (Wang et al. 2024c, d; Attia et al. 2024). The emergence of large multimodal models has driven further progress, particularly for tasks that require the alignment of speech and language. Models like Seed-ASR (Bai et al. 2024), Qwen2-Audio (Chu et al. 2024), and FireRedASR (Xu et al. 2025) offer greater flexibility and precision across diverse speech-related tasks.
@@ A.2 Context Compression
Recent approaches to context compression in large language models (LLMs) focus on improving the efficiency of handling long contexts while retaining important information. Chevalier et al. (2023) compresses long contexts into summary vectors, which are then used as soft prompts to improve model performance without significantly increasing the computational cost. Ge et al. (2023) utilizes a pre-trained autoencoder to generate compact memory slots for context. Li et al. (2023) proposes a method that prunes redundant information in the input context, effectively reducing its size while maintaining key contextual elements. LongLLMLingua (Jiang et al. 2023a, b) focuses on prompt compression for long-context scenarios, using a mechanism to enhance the LLM’s ability to retain key information in the compressed prompt. Cao et al. (2024) uses queries to guide the compression process, ensuring important contextual information is retained during compression even with high compression ratios.
@@ A.3 Contextualized ASR
Recent advances in Contextualized ASR have shown significant improvements by incorporating contextual information. A common approach is deep biasing end-to-end models. Huang et al. (2023) utilizes textual context phrases to assist in training ASR models, improving performance in recognizing longer, context-dependent sequences. Sudo et al. (2024) uses an editable bias phrase list to inject domain-specific knowledge into the ASR system. Yang et al. (2024c) incorporates text prompts into end-to-end ASR systems, enabling flexible transcription control and contextual understanding. Jayanthi et al. (2023) explores a two-stage retrieve-and-copy pipeline to retrieve related textual content based on similarity metrics. The other approaches utilize SpeechLLM models for enhanced contextual modeling capabilities. Lakomkin et al. (2024) uses LLaMA for speech recognition, treating the task as a mixed-modal language modeling problem and effectively incorporating textual contexts alongside audio inputs. Yang et al. (2024b) utilizes the WavLM Large model as a speech encoder, pre-trained on extensive audio data, and Vicuna as its LLM decoder.
@@ C.1 Dataset
We conduct experiments on the SlideSpeech (Wang et al. 2024b) dataset. This large-scale multimodal collection contains 1,705 YouTube-sourced conference videos, featuring synchronized 720p video streams, 16kHz audio recordings, and slide-derived OCR texts. The corpus uniquely integrates pre-processed slide content including extracted keywords and layout annotations. The training set is stratified into two subsets: L95 and S95, comprising 473 hours and 161 hours of audio data, respectively.
@@
To more fairly and comprehensively evaluate the performance of our proposed SAP2, we conduct experiments on the LibriSpeech (Panayotov et al. 2015) corpus. For the training and validation sets, we follow the approach in Le et al. (2021) to dynamically construct biasing lists. Specifically, for each sample, we first identify words in its ASR label that fall outside the common5k vocabulary. Then, we add $N$ (randomly selected between 400 and 800) distractors randomly sampled from the pool of 209.2k rare words. The test set remains consistent with Le et al. (2021). In line with most previous works, we evaluate our method under three settings: $N=\{100,500,1000\}$. We train separately on the 460-hour ”train-clean” set and the 500-hour ”train-other” set, and perform inference on the ”test-clean” and ”test-other” sets, respectively.